# DAASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples

### v1

@@ 4.2 Design Details
CIFAR-100 Cui et al. (2024) DI-FGSM 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
CW 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
PI-FGSM++ 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
AutoAttack 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
DiffPGD 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
DiffAttack 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
AdvAD 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
@@
Wang et al. (2023) DI-FGSM 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
CW 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
PI-FGSM++ 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
AutoAttack 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
DiffPGD 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
DiffAttack 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
AdvAD 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
@@
(Addepalli et al.) (2022) DI-FGSM 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
CW 100.00 87.21 96.48 77.93 60.74 84.47 95.56 0.0113 26.94
PI-FGSM++ 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
AutoAttack 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
DiffPGD 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
DiffAttack 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
AdvAD 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
@@
CIFAR-10 (Bartoldson et al.) (2024) DI-FGSM 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
CW 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
PI-FGSM++ 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
AutoAttack 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
DiffPGD 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
DiffAttack 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
AdvAD 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
@@
Cui et al. (2024) DI-FGSM 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
CW 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
PI-FGSM++ 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
AutoAttack 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
DiffPGD 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
DiffAttack 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
AdvAD 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
@@
Wang et al. (2023) DI-FGSM 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
CW 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
PI-FGSM++ 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
AutoAttack 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
DiffPGD 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
DiffAttack 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
AdvAD 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
@@
ImageNet-1k Salman et al. (2020) DI-FGSM 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
CW 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
PI-FGSM++ 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
AutoAttack 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
DiffPGD 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
DiffAttack 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
AdvAD 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
@@
5 Experiments and Results
Next we discuss our experimental setup and evaluations. All the reported results are obtained using NVIDIA A100 GPU.
5.1 Experimental Setup
5.1.1 Datasets and Models
Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 Krizhevsky et al. (2014), CIFAR-100 Krizhevsky (2009), and ImageNet Deng et al. (2009). For each dataset, we randomly sample 1,000 test images that are correctly classified by the victim model. Since standard models (without any defenses) are easily broken by existing attacks, we focus our evaluation on robust models curated by RobustBench Croce et al. (2020), a popular benchmark for adversarial robustness.
@@ 5.1.2 Post-Processing Defenses
To further test the limits of DAASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models). The post-processing defenses include JPEG compression Guo et al. (2017), Total Variation Minimization (TVM) Wang et al. (2018), Neural Representation Purifier (NRP) Naseer et al. (2020), and an ensemble of four methods: bit-depth reduction Xu (2018), JPEG, TVM, and Non-Local Means Denoising (NLM) Buades et al. (2011).
@@ 5.1.3 Evaluation Metrics
Figure 3: Comparison of DAASH performance before and after training, illustrating ASR and SSIM across different defense settings. “Trained DAASH" significantly outperforms the randomly weighted version (“Random DAASH").
5.1.4 Base Attacks and DAASH Training Details
We include 10 diverse base attacks in our attack pool: FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), CW Carlini and Wagner (2017), FAB Croce and Hein (2020a), TI-FGSM Dong (2019), NI-FGSM Lin et al. (2019), MI-FGSM Dong et al. (2017), DI-FGSM Xie et al. (2018), BIM Kurakin et al. (2016), and PI-FGSM++ Gao et al. (2020). Each attack offers unique advantages and covers a broad range of perturbation strategies. In addition, we include a None operation at each stage, which simply returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets. We run Algorithm 1 for $T=100$ epochs using the Adam optimizer with a learning rate of 0.001. After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example.
Figure 4: DAASH training progression over epochs.
5.2 Effectiveness of DAASH against Robust Models
Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DAASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet. As shown in the table, DAASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets. This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints. For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$. In contrast, DAASH achieves an ASR close to $100\%$. Moreover, DAASH yields better imperceptibility scores compared to these dedicated perceptual attacks. Similar trends are observed on CIFAR-10 and ImageNet: DAASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average. Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment.
Table 2: Effectiveness of DAASH with varying number of stages for the robust CIFAR-100 model – Cui et al. (2024).
@@ 5.2 Effectiveness of DAASH against Robust Models
Table 3: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model transferability.
@@
(Cui et al.) (2024) (Wang et al.) (2023) (Addepalli et al.) (2022) (Bartoldson et al.) (2024) (Cui et al.) (2024) (Wang et al.) (2023) (Salman et al.) (2020)
@@
Although DAASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower. Intuitively, since CW is one of the base attacks in DAASH, it could be expected to receive higher attention weights during optimization. However, this is not always the case. The reason lies in the mismatch between the objectives: CW uses a hinge-like loss that stops contributing gradients once misclassification is achieved, whereas our meta-loss continues to penalize the confidence of the true class beyond the decision boundary. As a result, after initial misclassification, CW receives lower weights in DAASH, which shifts focus to other attacks that still produce informative gradients. This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DAASH against additional defenses such as post-processing but comes at the cost of slightly reduced perceptual similarity.
5.3 Ablation Studies on DAASH
@@
Our DAASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima. In this section, we conduct ablation studies to demonstrate the importance of each component.
@@
Importance of Learnable Weights. To evaluate the benefit of using learnable weights, we compare the performance of DAASH against a variant where the weights for base attacks are initialized randomly and left untrained. Specifically, the random weights are sampled uniformly in the range $[0,1]$ and used directly without optimization. We repeat this random setup 100 times and report the average results.
@@
Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (measured by SSIM) across all types of robust models. Furthermore, Figure 4 illustrates the ASR and SSIM trends as training progresses over epochs. Both metrics exhibit clear improvement—ASR increases across all settings, and SSIM improves almost monotonically—highlighting the effectiveness of our meta-learning process. These results confirm that the learnable weighting mechanism in DAASH plays a crucial role.
@@
Importance of Multiple Stages. To assess the value of chaining multiple stages, we evaluate DAASH with varying numbers of stages. Table 2 reports the ASR for different stage configurations. The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that two stages are sufficient to escape bad local optima and refine the adversarial example generation. Adding more stages yields marginal improvements, indicating diminishing returns beyond two stages.
@@
Effectiveness against Unseen Defenses. We also examine whether the learned weights in DAASH are transferable across models and datasets. Table 3 shows that weights trained on one model or dataset can be applied to unseen models or datasets while still achieving near-optimal results. The variation in average ASR remains within 3%, and SSIM varies by less than 2. These results demonstrate that DAASH can generalize well across different settings/models.

### v2

@@ 5.1.1 Datasets and Robust Models
CIFAR-100 Cui2024 [cui2024decoupled] DI-FGSM 0.17 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
CW 0.11 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
PI-FGSM++ 0.22 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
AutoAttack 0.33 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
DiffPGD 4.21 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
DiffAttack 12.44 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
AdvAD 0.51 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
@@
Wang2023 [wang2023better] DI-FGSM 0.17 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
CW 0.12 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
PI-FGSM++ 0.22 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
AutoAttack 0.39 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
DiffPGD 4.77 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
DiffAttack 12.96 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
AdvAD 0.51 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
@@
Addepalli2022 [addepalli2022efficient] DI-FGSM 0.18 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
CW 0.13 100.00 86.67 96.45 79.04 58.72 84.18 94.94 0.0118 27.54
PI-FGSM++ 0.23 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
AutoAttack 0.39 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
DiffPGD 4.27 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
DiffAttack 12.86 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
AdvAD 0.54 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
@@
CIFAR-10 Bartold2024 [bartoldson2024adversarial] DI-FGSM 0.72 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
CW 0.62 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
PI-FGSM++ 0.83 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
AutoAttack 2.17 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
DiffPGD 4.80 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
DiffAttack 15.91 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
AdvAD 3.99 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
@@
Cui2024 [cui2024decoupled] DI-FGSM 0.16 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
CW 0.11 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
PI-FGSM++ 0.21 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
AutoAttack 0.41 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
DiffPGD 5.30 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
DiffAttack 12.88 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
AdvAD 0.54 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
@@
Wang2023 [wang2023better] DI-FGSM 0.17 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
CW 0.12 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
PI-FGSM++ 0.22 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
AutoAttack 0.41 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
DiffPGD 5.13 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
DiffAttack 13.49 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
AdvAD 0.54 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
@@
ImageNet-1k Salman2020 [salman2020adversarially] DI-FGSM 1.22 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
CW 0.58 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
PI-FGSM++ 1.75 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
AutoAttack 2.22 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
DiffPGD 10.40 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
DiffAttack 20.33 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
AdvAD 5.61 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
@@ 5.1.2 Post-Processing Defenses
To further test the limits of DASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models). The post-processing defenses include JPEG compression [guo2017countering], Total Variation Minimization (TVM) [wang2018adversarial], Neural Representation Purifier (NRP) [naseer2020self], and an ensemble of four methods: bit-depth reduction [xu2018feature], JPEG, TVM, and Non-Local Means Denoising (NLM) [buades2011non].
Table 2: Transferability Comparison on CIFAR-100 Dataset. S represents Surrogate models and T represents Target models.
S (↓) T (→) Attacks [cui2024decoupled] [wang2023better] [addepalli2022efficient] Avg.(↑) SSIM (↑)
[…]
CW 37.32 $73 \cellcolorgray$!1584.18 53.08 94.94
[…]
@@ 5.1.3 Evaluation Metrics
Figure 3: Comparison of DASH performance before and after training, illustrating ASR and SSIM across different defense settings. “Trained DASH” significantly outperforms the randomly weighted version (“Random DASH”).
5.1.4 Base Attacks and DASH Training Details
We include 10 diverse base attacks in our attack pool: FGSM [goodfellow2015explaining], PGD [madry2018towards], CW [carlini2017towards], FAB [croce2020minimally], TI-FGSM [dong2019evading], NI-FGSM [lin2019nesterov], MI-FGSM [dong2017boosting], DI-FGSM [xie2018improving], BIM [kurakin2016adversarial], and PI-FGSM++ [gao2020patch]. Each attack offers unique advantages and covers a broad range of perturbation strategies. For $\ell_{\infty}$-norm attacks, we set the perturbation budget to $\epsilon=\frac{7}{255}$. For different variants of FGSM evaluated in this paper, we use a momentum factor of 1. In TI-FGSM [dong2019evading] and DI-FGSM [xie2018improving], we set the resize_rate to 0.9, diversity_prob to 0.7, and random_start to False. In PI-FGSM++ [gao2020patch], we set the probability of using diverse inputs to 0.7 and the project_factor to 0.8. For the Fast Adaptive Boundary (FAB) attack [croce2020minimally], we set $\alpha_{\text{max}}=0.1$, overshooting factor $\eta=1.05$, and the backward step parameter $\beta=0.9$. For BIM [kurakin2016adversarial] and PGD [madry2017towards], we use a step size of $\alpha=0.01$ and number of iterations $T=10$. For the CW attack [carlini2017towards], we set the confidence parameter $c=10$, margin parameter $\kappa=0$, and optimized for 100 steps using a learning rate of $0.01$. In addition, we include a None operation at each stage, which returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
Table 3: Effectiveness of DASH with varying numbers of stages for the robust CIFAR-100 Cui2024 [cui2024decoupled] model.
@@ 5.1.4 Base Attacks and DASH Training Details
For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets. We run Algorithm 1 for $T=100$ epochs using the Adam [kingma2014adam] optimizer with a learning rate of 0.01. After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example. A visualization of this attack combination has been added in the supplimentary material.
Figure 4: DASH training progression over epochs.
5.2 Effectiveness of DASH against Robust Models
Table 4: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model weight transferability.
@@ 5.2 Effectiveness of DASH against Robust Models
Cui2024 [cui2024decoupled] Wang2023 [wang2023better] Addepalli22 [addepalli2022efficient] Bartold2024 [bartoldson2024adversarial] Cui2024 [cui2024decoupled] Wang2023 [wang2023better] Salman20 [salman2020adversarially]
@@
Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet. As shown in Table 1, DASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets. This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints. For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$. In contrast, DASH achieves an ASR close to $100\%$. Moreover, DASH yields better imperceptibility scores compared to these dedicated perceptual attacks. Similar trends are observed on CIFAR-10 and ImageNet: DASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average. Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment. We have obtained results with different parameter configurations for each compared attacks and reported the best results.
@@
Although DASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower. Intuitively, since CW is one of the base attacks in DASH, it could be expected to receive higher attention weights during optimization. However, this is not always the case. The reason lies in the mismatch between the objectives: CW uses a hinge-like loss that stops contributing gradients once misclassification is achieved, whereas our meta-loss continues to penalize the confidence of the true class beyond the decision boundary. As a result, after initial misclassification, CW receives lower weights in DASH, which shifts focus to other attacks that still produce informative gradients. This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DASH against additional defenses such as post-processing where CW fails.
5.3 Effectiveness of DASH on Black-Box Settings
Table 2 reports the attack performances of DASH in black-box scenarios. Additional preliminary results of using black-box attacks (e.g. Sign Flip [chen2020boosting], Rays [chen2020rays], Square [andriushchenko2020square], Sign Hunter [al2020sign]) as DASH base attacks can be found in the supplementary document (appendix).
Adversarial Examples from Surrogate Model. Table 2 shows that the adversarial examples generated with DASH for one model also perform well for other models. DASH achieves the strongest transferability, achieving the top average ASR for each model while also maintaining the best perceptual quality (SSIM). For Cui2024 [cui2024decoupled] and Wang2023 [wang2023better] models DASH outperforms the best baseline Diffattack with 14% and 16% improvement in average ASR, keeping the best SSIM score. For Addepalli2022 [addepalli2022efficient] DASH outperforms the best baseline AdvAD with almost a 5% improvement in average ASR, keeping the best SSIM score. This indicates that DASH consistently achieves good result in both transferability and imperceptibility scenarios.
5.4 Ablation Studies on DASH
@@
Our DASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima. In this section, we conduct ablation studies to demonstrate the importance of each component.
@@
Importance of Learnable Weights. To evaluate the benefit of using learnable weights, we compare the performance of DASH against a variant where the weights for base attacks are initialized randomly and left untrained. Specifically, the random weights are sampled uniformly in the range $[0,1]$ and used directly without optimization. We repeat this random setup 100 times and report the average results.
@@
Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (SSIM) across all types of robust models. Furthermore, Figure 4 illustrates the ASR and SSIM trends as training progresses over epochs. Both metrics exhibit clear improvement: ASR increases across all settings, and SSIM improves almost monotonically, highlighting the effectiveness of our meta-learning process. These results confirm that the learnable weighting mechanism in DASH plays a crucial role.
@@
Importance of Multiple Stages. To assess the value of chaining multiple stages, we evaluate DASH with varying numbers of stages. Table 3 reports the ASR for different stage configurations. The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that multi-stage can help escape bad local optima and refine the adversarial example generation. Adding more stages may yield improvements.
@@
Effectiveness against Unseen Defenses. We also examine whether the learned weights in DASH are transferable across models and datasets. Table 4 shows that weights trained on one model or dataset can be applied to unseen models or datasets while still achieving near-optimal results. The variation in average ASR remains within 3%, and SSIM varies by less than 2. These results demonstrate that DASH can generalize well across different settings/models.